# Deep Ignorance: Filtering Pretraining Data Builds Tamper-Resistant Safeguards into Open-Weight LLMs

### v1

@@ 6.1 Implications for Open- and Closed-Weight Model Risk Management
The machine unlearning and adversarial robustness fields have long struggled to build durable tamper-resistant safeguards into LLMs (Huang et al., 2024; Qi et al., 2024b; Che et al., 2025). This may be due, in part, to a past focus on post-training techniques. While these methods are useful, post-training tools persistently struggle to deeply modify a model’s internal knowledge representations in a way that prevents fine-tuning-facilitated recovery (Jain et al., 2023). Here, by shifting our focus from post-training to pre-training interventions, we demonstrate state-of-the-art tamper-resistance for up to 10,000 steps and 300M tokens of adversarial fine-tuning. We also find useful synergies between training data filtering and existing post-training safeguards (Zou et al., 2024; Sheshadri et al., 2024).
We have shown that pretraining data filtering is a powerful tool for improving open-weight safety against adversarial and non-adversarial forms of tampering alike. Based on these results (Section 3.2), we argue that training data filtering can be a useful component of risk management strategies for open-weight LLMs. However, open-weight model risk management remains fundamentally challenging because the downstream risks of open-weight models depend, in part, on the resources and goals of external actors (which are outside the developer’s control). As a result, comprehensive risk management for open-weight models may require additional risk monitoring and mitigation strategies aside from what model-based safeguards can offer alone. Toward these ends, it will be useful to continue building the science of assessing, monitoring, and mitigating risks in the open-weight ecosystem.
@@ 6.2 Does pretraining data filtering help to mitigate all types of harmful behaviors?
Observing that harmful training data could sometimes lead to more safe models, Maini et al. (2025) and Li et al. (2025a) speculated that language models sometimes need to ‘understand’ harmful behaviors to be able to effectively resist exhibiting them. They hypothesized that “Safety is not about censorship,” (Maini et al., 2025) and that “Bad data may lead to good models,” (Li et al., 2025a). However, this does not fully explain successful results from here and Lee et al. (2025). Based on all of these findings, we speculate that this hypothesis only applies to emergent propensities (e.g., toxicity, attempted compliance with harmful requests, aligning with a particular set of principles) which do not require precice knowledge to be exhibited. However, we suspect that this hypothesis does not apply to knowledge (e.g., scientific- or engineering-relevant facts) which is precise in nature and arises only from a small subset of training documents. For suppressing unwanted knowledge, we hypothesize that filtering pretraining data can build robust safeguards into LLMs. Along these lines, future work to study the strengths and limitations of training data curation may help to better inform safeguards research.
@@ 6.3 Limitations
Our experiments here are limited to a particular set of models and experimental configurations. We only study unimodal 6.9B parameter language models without instruction fine-tuning. Due to logistical and financial considerations, we only trained these models on 550B tokens until they surpassed 50% performance on the standard WMDP-Bio eval. They do not have competitive capabilities with other open LLMs of similar sizes. Additionally, we also only experiment in the context of biothreat proxy knowledge. While we have done our best to tune and configure our attacks to be as strong as possible, adversarial evaluation approaches like ours can only offer a lower bound for the worst possible case behavior (Gal, 2024). It is likely that different tampering attacks could more effectively elicit unwanted information from our models than our implemented attacks.
@@ 6.5 Future Work
We would like to thank Yejin Choi, Liwei Jiang, Arthur Conmy, Grace Braithwaite, May Dixit, Kateryna Halstead, James Zhang, Aytunç Ilhan, Peter Gebauer, A. Feder Cooper, Adam Gleave, Geoffrey Irving, Pietro Lesci, Ian McKenzie, Samuel Ratnam, Paul Rottger, Lydia O’Brien, Cameron Tice, Blake Bullwinkel, Nora Belrose, and Aviya Skowron for helpful discussions. Alex Robey and Alexandra Souly also provided valuable methodological input. Jai Patel coordinated collaboration logistics between EleutherAI and UK AISI. Iman Syed offered support related to compute behind our tampering experiments. Kyle O’Brien was partially supported financially by the Cambridge ERA:AI Fellowship.
@@ D.1.3 Classifier Filter
Optimizer AdamW (Loshchilov & Hutter, 2017)
@@ D.2 Language Model Training Setup
We train models using the EleutherAI GPT-NeoX library (Andonian et al., 2023) on a cluster of 128 Nvidia H100s. Models follow an identical architecture to Pythia 6.9B (Biderman et al., 2023). We train with a sequence length of 2,048 tokens and an effective batch size of 4,194,304 tokens. We use the same tokenizer as GPT-NeoX (Black et al., 2022) and Pythia (Biderman et al., 2023).
@@ D.2.3 Hyperparameters
Activation Function GELU (Hendrycks & Gimpel, 2016)
@@ D.3.1 Step 1: Calculate Throughput (Observed H100 FLOPS)
We can calculate the Model FLOPs Utilization (MFU) to quantify training efficiency. Nvidia reports (Bekman, 2023-2024) the peak FLOPS for dense BF16 operations as 989e12. The following MFU of 0.56 suggests healthy efficiency:
@@
The above MFU calculation uses the 989e12 max provided by Nvidia. However, other sources have suggested that even this maximum is out of reach in synthetic replications. The Machine Learning Engineering Open Book (Bekman, 2023-2024) suggests that 794.5e12 is the Maximum Achievable Matmul FLOPS (MAMF) for H100s. Using MAMF, we arrive at:
@@ I.5 WMDP Cloze-Style Compatability
We use OpenAI o3 (OpenAI, 2025) with the following system prompt to classify which WMDP-Bio benchmark samples can be converted to cloze-style. The benchmark prompt is passed in as the user message.

### v2

@@ 6.1 Implications for Open- and Closed-Weight Model Risk Management
The machine unlearning and adversarial robustness fields have long struggled to build durable tamper-resistant safeguards into LLMs (Huang et al., 2024; Qi et al., 2024b; Che et al., 2025). This may be due, in part, to a past focus on post-training techniques. While these methods are useful, post-training tools persistently struggle to deeply modify a model’s internal knowledge representations in a way that prevents fine-tuning-facilitated recovery (Jain et al., 2023). Here, by shifting our focus from post-training to pre-training interventions, we demonstrate state-of-the-art tamper-resistance for up to 10,000 steps and 300M tokens of adversarial fine-tuning. We also find useful synergies between training data filtering and existing post-training safeguards (Zou et al., 2024; Sheshadri et al., 2024). Based on these results, we argue that training data filtering can be a useful component of risk management strategies for open-weight LLMs.
Is 10,000 steps and 300M tokens of tamper resistance enough? Here, we find that filtering pretraining data makes models more than an order of magnitude more resistant to tampering than state-of-the-art post-training baselines. But in absolute terms, 10,000 steps and 300M tokens of fine-tuning is not a very large amount. At the point at which an adversary could perform 100 steps of fine-tuning, they almost certainly have the compute, code, configuration, and expertise they need to do more. Running fine-tuning for a few thousand steps longer would incur a very small marginal cost. However, we believe that obtaining high-quality data for adversarial fine-tuning may become a key bottleneck. Here, we work with biothreat-proxy data using the WMDP-bio dataset, which consists of 24,453 documents on biothreat proxy subjects. Li et al. (2024b) reported that making the WMDP datasets and evaluations cost over \$200,000 USD. This suggests that collecting high-quality fine-tuning data may be a significant obstacle, costing potentially tens of thousands of dollars and significant expert input. Compared to WMDP-bio, which is a fairly harmless proxy constructed for research purposes, it may be substantially harder to obtain the expertise to construct genuinely info-hazardous datasets. However, the relationship between the effort required to construct infohazardous datasets and the effectiveness of fine-tuning on them is not well understood. We leave investigating this relationship to future work.
Overall, open-weight model risk management remains fundamentally challenging because the downstream risks of open-weight models depend, in part, on the resources and goals of external actors (which are outside the developer’s control). As a result, comprehensive risk management for open-weight models may require additional risk monitoring and mitigation strategies aside from what model-based safeguards can offer alone. Toward these ends, it will be useful to continue building the science of assessing, monitoring, and mitigating risks in the open-weight ecosystem.
@@ 6.2 Does pretraining data filtering help to mitigate all types of harmful behaviors?
Observing that harmful training data could sometimes lead to more safe models, Maini et al. (2025) and Li et al. (2025a) speculated that language models sometimes need to ‘understand’ harmful behaviors to be able to effectively resist exhibiting them. They hypothesized that “Safety is not about censorship,” (Maini et al., 2025) and that “Bad data may lead to good models,” (Li et al., 2025a). However, this does not fully explain successful results from here and Lee et al. (2025). Based on all of these findings, we speculate that this hypothesis only applies to emergent propensities (e.g., toxicity, attempted compliance with harmful requests, aligning with a particular set of principles) which do not require precise knowledge to be exhibited. However, we suspect that this hypothesis does not apply to knowledge (e.g., scientific- or engineering-relevant facts) which is precise in nature and arises only from a small subset of training documents. For suppressing unwanted knowledge, we hypothesize that filtering pretraining data can build robust safeguards into LLMs. Along these lines, future work to study the strengths and limitations of training data curation may help to better inform safeguards research.
@@ 6.3 Limitations
Our experiments here are limited to a particular set of models and experimental configurations. We only study unimodal 6.9B parameter language models without instruction fine-tuning. Due to logistical and financial considerations, we only trained these models on 550B tokens until they surpassed 50% performance on the standard WMDP-Bio eval. They do not have competitive capabilities with other open LLMs of similar sizes. We also only experiment in the context of biothreat proxy knowledge evaluated with multiple-choice questions (Khatun and Brown, 2024). While we have done our best to tune and configure our attacks to be as strong as possible, adversarial evaluation approaches like ours can only offer a lower bound for the worst possible case behavior (Gal, 2024). It is likely that different tampering attacks could more effectively elicit unwanted information from our models than our implemented attacks.
@@ 6.5 Future Work
We would like to thank Yejin Choi, Liwei Jiang, Arthur Conmy, Grace Braithwaite, May Dixit, Kateryna Halstead, James Zhang, Aytunç Ilhan, Peter Gebauer, A. Feder Cooper, Adam Gleave, Pietro Lesci, Ian McKenzie, Samuel Ratnam, Paul Rottger, Lydia O’Brien, Cameron Tice, Blake Bullwinkel, Nora Belrose, Patricia Paskov and Aviya Skowron for helpful discussions. Alex Robey and Alexandra Souly also provided valuable methodological input. Jai Patel coordinated collaboration logistics between EleutherAI and UK AISI. Iman Syed offered support related to compute behind our tampering experiments. Kyle O’Brien was partially supported financially by the Cambridge ERA:AI Fellowship.
@@ D.1.3 Classifier Filter
Optimizer AdamW (Loshchilov and Hutter, 2017)
@@ D.2 Language Model Training Setup
We train models using the EleutherAI GPT-NeoX library (Andonian et al., 2023) on a cluster of 128 Nvidia H100s. Models follow an identical architecture to Pythia 6.9B (Biderman et al., 2023). We train with a sequence length of 2,048 tokens and an effective batch size of 4,194,304 tokens. We use the same tokenizer as GPT-NeoX (Black et al., 2022) and Pythia (Biderman et al., 2023). All pre-training runs for this project were conducted under a compute contract totaling \$476k USD.
@@ D.2.3 Hyperparameters
Activation Function GELU (Hendrycks and Gimpel, 2016)
@@ D.3.1 Step 1: Calculate Throughput (Observed H100 FLOPS)
We can calculate the Model FLOPs Utilization (MFU) to quantify training efficiency. Nvidia reports (Bekman, 2023) the peak FLOPS for dense BF16 operations as 989e12. The following MFU of 0.56 suggests healthy efficiency:
@@
The above MFU calculation uses the 989e12 max provided by Nvidia. However, other sources have suggested that even this maximum is out of reach in synthetic replications. The Machine Learning Engineering Open Book (Bekman, 2023) suggests that 794.5e12 is the Maximum Achievable Matmul FLOPS (MAMF) for H100s. Using MAMF, we arrive at:
@@ I.5 WMDP Cloze-Style Compatability
We use OpenAI o3 (OpenAI, 2025d) with the following system prompt to classify which WMDP-Bio benchmark samples can be converted to cloze-style. The benchmark prompt is passed in as the user message.